OpenStudy (anonymous):

Can somebody please help me with finding limits. I have four hw questions and i just dont understand .

OpenStudy (anonymous):

1. lim f(x), f(x)={4-x x<-1} x--> -1 {5 x=-1} {x+6 x>-1}

OpenStudy (anonymous):

find the limit from right and left using the correct definitions , if were both equal then it exists , otherwise it doesn't exist

OpenStudy (anonymous):

could you explain further?

OpenStudy (anonymous):

@Ahmad1

OpenStudy (anonymous):

\[\lim_{x \rightarrow -1^-}f(x)=\lim_{x \rightarrow -1^-}4-x=5\] \[\lim_{x \rightarrow -1+}f(x) =\lim_{x \rightarrow -1+}6+x=5\] the limit from both sides equal to five then we say that: \[\lim_{x \rightarrow -1}f(x)=5\] further more, the value of the limit equaled f(-1), this means that the not only the does exist, also the function is contious on R